# Dynamic Attention-Guided Context Decoding for Mitigating Context Faithfulness Hallucinations in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) often suffer from context faithfulness hallucinations, where outputs deviate from retrieved information due to insufficient context utilization and high output uncertainty. Our uncertainty evaluation experiments reveal a strong correlation between high uncertainty and hallucinations. We hypothesize that attention mechanisms encode signals indicative of contextual utilization, validated through probing analysis. Based on these insights, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a lightweight framework that integrates attention distributions and uncertainty signals in a single-pass decoding process. Experiments across QA datasets demonstrate DAGCD’s effectiveness, achieving significant improvements in faithfulness and robustness while maintaining computational efficiency.
@@ 1 Introduction
Large Language Models (LLMs) Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023) excel at generating fluent and contextually relevant responses. However, they often struggle with producing factually accurate outputs, particularly when relying on external information Vu et al. (2023). Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Guu et al. (2020) mitigates this by grounding outputs in retrieved context, making it effective for tasks like question answering and reasoning Gao et al. (2023); Fan et al. (2024). However, models often fail to faithfully utilize retrieved context, resulting in context faithfulness hallucinations, where outputs deviate from the retrieved context Huang et al. (2023a); Ji et al. (2023). These issues undermine RAG systems’ reliability, particularly in high-stakes applications requiring factual accuracy Chuang et al. (2024a).
Figure 1: Analysis of generation distribution when concatenating context: ($a_{1},a_{2}$) the model’s uncertainty when generating correct versus wrong answers; ($b$) for wrong answers, the ranking of the gold answer token in the distribution, and the probability difference between the gold answer token and the Top-1 predicted token.
@@
Existing methods, such as CAD Shi et al. (2024b) and COIECD Yuan et al. (2024), attempt to mitigate context faithfulness hallucinations by dynamically adjusting decoding distributions through generation distribution comparisons or token-level uncertainty signals. While effective to some extent, these methods face several key limitations: limited interpretability, degraded performance when context-agnostic and context-aware outputs differ significantly, and high computational overhead due to multiple decoding passes. These issues highlight the need for lightweight and interpretable approaches that can effectively utilize retrieved context while maintaining scalability.
To further explore these issues, we analyze the relationship between output uncertainty and context faithfulness hallucinations. As shown in Figure 1, we observe a strong correlation between context hallucinations and high entropy in the generation distribution. Specifically, incorrect generations often reflect significant uncertainty through high entropy. Notably, in cases of wrong answers, context tokens corresponding to the correct answer often appear within the top-10 ranks but are assigned insufficient confidence. This indicates that the model’s internal mechanisms encode interpretable signals to indicate context utilization.
Building on our analysis and findings in interpretability studies, we hypothesize that attention mechanisms in Transformer models encode signals indicative of contextual utilization. Attention modules play a critical role in aligning generated tokens with their corresponding context tokens by distributing focus across input tokens Geva et al. (2023); Jin et al. (2024b). To validate our hypothesis, we trained a probing classifier using Logistic Regression on attention distributions, achieving over 97% accuracy in distinguishing contextually relevant tokens. These results demonstrate that attention weights can serve as a lightweight and interpretable proxy for assessing the effective incorporation of retrieved evidence during generation.
@@
Motivated by these findings, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a novel algorithm to mitigate context faithfulness hallucinations. Inspired by the copy-generator framework See et al. (2017); Xu et al. (2020b), DAGCD integrates attention weights to estimate the relevance of tokens in the retrieved context, dynamically adjusting output probabilities. Entropy-based uncertainty measures further refine these adjustments, amplifying underconfident but contextually relevant tokens. By unifying these strategies, DAGCD aligns outputs with context while maintaining computational efficiency.
@@
    Comprehensive analysis of context faithfulness hallucinations: We reveal a strong correlation between context faithfulness hallucinations and output uncertainty through entropy-based analysis, revealing the model’s challenges in aligning with retrieved context under high uncertainty.
@@
    Lightweight and efficient decoding: DAGCD operates in a single decoding pass, integrating attention signals and uncertainty measures without additional overhead, achieving improved decoding efficiency.
@@
    Extensive validation across datasets and models: DAGCD outperforms greedy decoding across multiple QA datasets, achieving an average F1 score improvement of 8.95% on pretrained models and 1.91% on instruction-tuned models, demonstrating its robustness and scalability.
@@ 2 Why Can’t Generate Faithful Answers?
LLMs often fail to generate responses that are faithful to the concatenated context. We analyzed (1) the impact of context concatenation on output uncertainty (§2.1) and (2) the ranking of contextual answer tokens within predictions (§2.2). Our findings indicate that high output uncertainty and low confidence in contextually relevant tokens are critical factors contributing to this issue.
@@ 2.1 Uncertainty Leads to Unfaithful Answers
To assess the relationship between uncertainty and response faithfulness, we focus on two widely used metrics: entropy, which measures the overall uncertainty in the output distribution Xu et al. (2020a); van der Poel et al. (2022); Huang et al. (2023b), and top-1 token probability, which reflects the model’s confidence in its highest-ranked prediction Yao et al. (2023); Luo et al. (2024). Higher entropy indicates greater uncertainty, while higher top-1 probability corresponds to greater confidence in the model’s predictions.
Experimental Setup We compute the normalized entropy and top-1 token probability for the first generated token. Normalized entropy is defined as:

### v2

@@ Abstract
Large language models (LLMs) often exhibit Context Faithfulness Hallucinations, where outputs deviate from retrieved information due to incomplete context integration. Our analysis reveals a strong correlation between token-level uncertainty and hallucinations. We hypothesize that attention mechanisms inherently encode context utilization signals, supported by probing analysis. Based on these insights, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a lightweight framework that leverages attention distributions and uncertainty signals in a single-pass decoding. Experiments on open-book QA datasets demonstrate DAGCD’s effectiveness, yielding significant improvements in faithfulness and robustness while preserving computational efficiency. [^1] Our code is available at uestc-huangyw/DAGCD.
@@ 1 Introduction
Large Language Models (LLMs) Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023) excel in generating fluent and contextually relevant responses. However, they often struggle with factual accuracy, especially when relying on external information. Vu et al. (2023). Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Guu et al. (2020) mitigates this by grounding outputs in retrieved context, making it effective for tasks like question answering and reasoning Gao et al. (2023); Fan et al. (2024). However, models often fail to faithfully utilize retrieved context, resulting in Context Faithfulness Hallucinations, where outputs deviate from the retrieved context Huang et al. (2023a); Ji et al. (2023).
@@
These hallucinations undermine the reliability of RAG systems, particularly in critical domains where factual accuracy is paramountChuang et al. (2024a). Existing methods, such as CAD Shi et al. (2024b) and COIECD Yuan et al. (2024), attempt to mitigate context faithfulness hallucinations by dynamically adjusting decoding distributions through token-level probability distribution comparisons or token-level uncertainty signals. While effective to some extent, these methods face several key limitations: limited interpretability, degraded performance when context-agnostic and context-aware outputs differ significantly, and computational complexity due to multiple decoding passes.
Figure 1: Analysis of the token-level probability distribution after context concatenation: (a1, a2) the model’s uncertainty when generating correct versus wrong answers, measured by NE and MSP; (b) for wrong answers, the ranking of the golden answer token within the token-level probability distribution.
To better understand context faithfulness hallucinations and explore potential solutions, we take an internally-driven approach, analyzing intrinsic model signals that may explain why retrieval-augmented large language models struggle to utilize retrieved context faithfully Liang et al. (2024). Motivated by prior research linking token-level uncertainty to factual hallucinations Chuang et al. (2024b); Das et al. (2025), we examine whether token-level probability distribution entropy correlates with context faithfulness hallucinations.
Our analysis reveals a strong correlation between higher uncertainty and context faithfulness hallucinations. Specifically, as shown in Figure 1 (a1), wrong answers exhibit higher entropy in the token-level probability distribution, indicating greater uncertainty in generation. Even for correct answers (Figure 1 (a2)), it often assigns low confidence to the highest-ranked token, suggesting incomplete integration of retrieved context. Notably, in incorrect responses, most gold answer tokens appear in the top 10 in the token-level probability distribution, but are not assigned the highest probability (Figure 1 (b)), implying that the model identifies relevant context but fails to prioritize it effectively.
These findings indicate that while models retrieve relevant context, they struggle to integrate and prioritize it during generation. Since most gold answer tokens appear within the top 10, an effective strategy to mitigate context faithfulness hallucinations is to dynamically identify and prioritize these tokens during generation. This requires detecting reliable signals that indicate how retrieved context influences the model’s predictions. Attention mechanisms in Transformer models naturally emerge as a key source of such signals, since they facilitate information flow between tokens Olsson et al. (2022a); Meng et al. (2022a); Geva et al. (2023). We hypothesize that attention distributions encode intrinsic indicators of context utilization.
To validate our hypothesis, we trained a probing classifier using Logistic Regression on attention distributions, achieving over 0.99 AUC in distinguishing contextually relevant tokens. Even with just 100 training samples, the classifier demonstrated strong generalization across in-domain and cross-domain test sets, indicating that attention distributions inherently encode context utilization signals. These results support attention-based context utilization as a fundamental mechanism in LLMs. By leveraging these intrinsic signals, attention distributions provide a lightweight and interpretable means to assess how models integrate retrieved context into their predictions.
@@
Motivated by these findings, we propose Dynamic Attention-Guided Context Decoding (DAGCD), a novel method to mitigate context faithfulness hallucinations. Inspired by the copy-generator framework See et al. (2017); Xu et al. (2020), DAGCD integrates attention weights to estimate the relevance of tokens in the retrieved context, dynamically adjusting output probabilities. Additionally, token-level uncertainty guides these adjustments by emphasizing underconfident yet contextually relevant tokens. By combining these strategies, DAGCD ensures output alignment with the retrieval context and maintains efficiency.
@@
    Comprehensive analysis of context faithfulness hallucinations: We identify a strong correlation between token-level uncertainty and context faithfulness hallucinations, showing that incorrect responses exhibit higher entropy and retrieved context is often recognized but not effectively prioritized.
@@
    Lightweight and efficient decoding: DAGCD operates in a single decoding pass, integrating attention signals and uncertainty measures without additional overhead, improving efficiency.
@@
    Extensive validation across datasets and models: DAGCD outperforms greedy decoding across multiple QA datasets, improving EM by 17.67% on pretrained models and 2.25% on instruction-tuned models, demonstrating robustness and scalability.
@@ 2 Why Can’t Generate Faithful Answers?
Token-level uncertainty is closely related to factual hallucinations, as models often exhibit higher entropy in the token-level probability distribution when generating factually incorrect outputs Chuang et al. (2024b); Das et al. (2025). While uncertainty measures help detect hallucination-prone predictions, most studies focus on factual hallucinations, where responses are incorrect without retrieved context. In contrast, context faithfulness hallucinations arise when models rely on retrieved information but generate misaligned or contradictory outputs. Despite their significance in RAG, their relationship with uncertainty remains unclear. Inspired by prior research, we investigate whether unfaithful responses in RAG exhibit higher entropy and whether contextually relevant tokens are recognized but assigned insufficient confidence.
@@ 2.1 Uncertainty Leads to Unfaithful Answers
To assess the relationship between uncertainty and response faithfulness, we use two common metrics: Normalized Entropy (NE), which measures the overall uncertainty in the token-level probability distribution Huang et al. (2023b), and Maximum Softmax Probability (MSP), which quantifies the model’s confidence in the highest probability within the token-level probability distribution Hendrycks and Gimpel (2017). Higher NE indicates greater uncertainty, while higher MSP corresponds to greater confidence in the predictions of the model. For detailed experimental descriptions, see Appendix A.
Results and Analysis
Figure 1 (a1) and (a2) illustrates the Normalized Entropy and Maximum Softmax Probability of the token-level probability distribution when the model produces correct and wrong answers. The model exhibits higher uncertainty for wrong answers, with an average NE of 0.36 compared to 0.29 for correct cases, and a lower average MSP of 0.25 compared to 0.41. Notably, there is a substantial overlap between the correct (blue) and wrong (red) cases in the figure, indicating that even correct answers often exhibit high uncertainty. We also analyzed the correlation between prediction accuracy and uncertainty, and the results demonstrate a significant negative correlation, further confirming that token-level uncertainty is strongly associated with unfaithful answers. For detailed results see Appendix A.4.
2.2 LLM is Actually Utilizing Context
Our previous analysis links token-level uncertainty to unfaithful responses, showing that incorrect outputs often have higher entropy and lower confidence. However, this does not mean the model ignores retrieved context. A key question remains: Do incorrect responses imply that LLMs have failed to leverage the retrieved context?
To investigate this, we examine the ranking of gold answer tokens in the token-level probability distribution for incorrect responses. If these tokens frequently rank high but are not assigned the highest probability, it suggests the model identifies relevant context but fails to prioritize it effectively.
Experimental Setup
We analyze incorrect responses from §2.1 by examining the ranking distribution of gold answer tokens in the token-level probability distribution of the first generated token after context concatenation. The ranks are grouped into intervals, and we compute the proportion of gold answer tokens within each rank interval.
Results and Analysis
As shown in Figure 1 (subplot b), when the model generates an incorrect response, 43% of cases have the gold answer token ranked 2nd to 4th based on token probabilities, and 66% have it ranked within the top 10 in the token-level probability distribution. Moreover, as shown in Figure 6, the average probability gap between the gold answer token and the highest-probability token remains relatively small: 0.14 for ranks between 2 and 4, and 0.24 for ranks beyond 30.
These findings indicate that in context faithfulness hallucination scenarios, the model recognizes relevant context tokens but fails to prioritize them effectively, limiting their impact on the generated output. This highlights an incomplete integration of retrieved context, emphasizing the need for improved context incorporation strategies to enhance faithfulness and mitigate hallucinations.
3 Context Utilization Signal in Attention
Our analysis in Section 2 highlights that while models often identify relevant context tokens, they fail to assign them sufficient confidence, leading to context faithfulness hallucinations. To better understand this phenomenon, we seek reliable signals that indicate which retrieved context tokens are effectively utilized by the model during generation.
Due to their role in integrating and propagating information across tokens, attention mechanisms naturally emerge as a key candidate for capturing such signals Olsson et al. (2022a); Meng et al. (2022a); Geva et al. (2023). We hypothesize that attention distributions encode intrinsic indicators of context utilization, providing a lightweight and interpretable means to assess how models incorporate retrieved context into their outputs.
3.1 Attention Ratio
A key challenge in analyzing attention weights is the noise from non-context tokens (e.g., delimiters) caused by attention sink effects Bondarenko et al. (2021); Xiao et al. (2024). Additionally, attention magnitudes vary across heads and layers, complicating feature comparison Jain and Wallace (2019); Vig and Belinkov (2019). To address these issues, we introduce the attention ratio, a normalized measure that captures how much attention a retrieved context token receives relative to the total attention assigned within the context. For a given token $j$ in the retrieved context $C$, the attention ratio at the $h$-th head in the $l$-th layer is defined as:
$r_{l,h}^{j}=\frac{a_{l,h}^{j}}{\sum_{j\in C}a_{l,h}^{j}}$ (1)
where $a_{l,h}^{j}$ denotes the raw attention weight assigned to token $j$. This ratio quantifies the relative importance of $j$ within the retrieved context for a specific attention head.
To construct token-level features, we aggregate attention ratios across all heads:
This feature vector represents the distribution of attention across layers and heads, enabling a structured assessment of context token importance.
Figure 2: Cross Domain Validation. One dataset as the training set (X-axis) and the remaining datasets as the test sets, showing the AUC on the test sets (Y-axis).
3.2 Experimental Setup
To validate our hypothesis and investigate whether this mechanism generalizes across different datasets, training sizes, and prompt templates, we conduct the following experiments.
Data Construction We constructed the dataset by randomly selecting samples from the MrQA training set dataset (contains six open-book QA datasets in different domains) Fisch et al. (2019), focusing on cases where the model’s output changed from incorrect to correct after context concatenation Meng et al. (2022b). These cases indicate that the model successfully leveraged the retrieved context to produce the correct answer. Context tokens were labeled as positive (utilized) if they corresponded to the gold answer, and negative (non-utilized) otherwise. Using these labels, we extracted attention ratio feature vectors $\mathbf{v}_{j}$ to train a Logistic Regression (LR) classifier.
3.3 Results
Cross Domain Validation
We tested the classifier on six sub-datasets from different domains. Specifically, we selected one dataset as the training set and tested the performance on the remaining datasets (each sub-dataset construct 500 samples, contains 250 positive and 250 negative samples).
As shown in Figure 2, the classifier achieves an average AUC above 0.99 across all datasets and LLMs. The results indicate that the context utilization signal in attention is data-independent.
Training Data Size Validation
Building on the cross domain experiment, we further tested the impact of training set size on classifier performance. Specifically, we trained the model using data constructed from one sub-dataset and evaluated its performance on the remaining sub-datasets.
Figure 3 shows the variation in model AUC with different training data sizes (where "Train size = 100" refers to a training set constructed with 50 positive and 50 negative samples). The results indicate that even with only 100 samples, the model’s AUC exceeds 0.96, and the performance improvement becomes limited as the data size increases. This demonstrates that the LR classifier trained using the attention ratio exhibits strong data-efficiency.
Figure 3: Training Data Size Validation. Training sets of varying sizes were constructed from a single dataset (HotpotQA), and evaluate on remain datasets.
Additional Results
We then examined the impact of different prompt templates, and the results indicate that the classifier consistently maintains high performance regardless of the prompt template used. For Details, see Appendix B.
To better apply the classifier to practical tasks, we conducted a detailed analysis of the importance of different features and the relationships between them. The results show that the classifier using the top-K features outperforms the one using the full feature set. Furthermore, the attention heads exhibit Concentration and Complementarity features. For Details, see Appendix C.
Conclusion: A Fundamental Mechanism in Transformer-based LLMs
The consistent generalization of attention-based context utilization across datasets, data sizes, and prompts reinforces its role as a fundamental mechanism in LLMs. Our findings show that attention heads encode robust context integration signals, providing a lightweight and interpretable way to assess how models incorporate retrieved context.
4 Method
Inspired by the copy-generation mechanism See et al. (2017); Xu et al. (2020), we propose Dynamic Attention-Guided Context Decoding (DAGCD) to mitigate context faithfulness hallucinations. DAGCD leverages utilization signals to dynamically guide the generation process, focusing on relevant contextual tokens. It integrates three steps: detecting utilized context tokens during inference (§4.1), constructing a utilization distribution (§4.2), and adjusting the token-level probability distribution to enhance contextual utilization (§4.3).
4.1 Context Utilization Detection
Context Utilization Detector To identify contextually relevant tokens during inference, we utilize a Logistic Regression (LR) classifier trained on attention-based utilization signals. Our analysis in Section 3.3 shows that selecting the most informative attention heads improves generalization. Thus, we construct the Context Utilization Detector based on the top-K most important attention heads.
Feature Data Collection To obtain feature vectors, we extract attention distributions at the current decoding step, as illustrated in Figure 4. Specifically, we take the last row of the attention map for each selected head $h_{k}\in H$ ($H$ is the set of top-K attention heads). To ensure focus on relevant information, we filter out non-contextual tokens, such as query tokens and placeholder tokens in templates.
The top-K feature vector for each contextual token $j$ is then constructed as:
$\mathbf{v}_{j}^{(K)}=[r_{h_{1}}^{j},r_{h_{2}}^{j},\dots,r_{h_{K}}^{j}]$ (3)
where $r_{h_{k}}^{j}$ represents the normalized attention ratio of token $j$ in attention head $h_{k}$.
4.2 Utilization Distribution Construction
The context utilization detector identifies which tokens are utilized but does not quantify the degree of utilization for each token. To address this, we compute a utilization score $s_{j}$ for each token $j$ by aggregating attention ratios from selected attention heads, weighted by their normalized feature coefficients $w_{k}$. Tokens classified as unused by the detector are directly assigned a score of zero.
$s_{j}=\sum_{k=1}^{K}(r_{h_{k}}^{j}\times w_{k}),\ w_{k}=\frac{c_{k}}{\sum_{k=1}^{K}c_{k}}$ (4)
where $r_{h_{k}}^{j}$ is the normalized attention ratio of token $j$ in attention head $h_{k}$, and $w_{k}$ is the importance weight assigned to head $h_{k}$. The coefficient $c_{k}$ is learned from the LR classifier, representing the contribution of each head to context utilization.
The utilization distribution $U$ represents a probability distribution over context tokens, normalized based on their utilization scores:
where $u_{i}$ denotes the utilization probability of token $i$, $N$ is the vocabulary size. Tokens either absent from the context or classified as non-utilized by the detector ($s_{i}=0$) are assigned $u_{i}=0$.
Top-Rank Constraint
To enhance the reliability of generation adjustments, our approach applies a top-rank restriction, ensuring modifications focus on plausible tokens. Specifically, we define $U_{\text{top}}$ as the subset of the utilization distribution $U$ corresponding to tokens ranked within the top-$R$ positions of the generation distribution.
This design builds on prior work Li et al. (2023); Chuang et al. (2024b), addressing context faithfulness hallucination challenges while leveraging our observation that correct context tokens usually appear within the top-ranked positions in the token-level probability distribution, even when the model generates incorrect answers. Through this constraint, we reduce the risk of amplifying irrelevant or nonsensical tokens, preserving output integrity.
4.3 Generating More Faithful Answers
DAGCD adjusts token probabilities based on token-level uncertainty, measured using the normalized entropy $H_{\text{norm}}(P)$ of the token-level probability distribution. High entropy indicates greater uncertainty and a higher risk of generating contextually inconsistent responses. Since entropy correlates with uncertainty but lacks a fixed numerical relationship, we introduce a scaling factor $\alpha$ to compensate for model-specific entropy variations.
Adjustments are applied only when utilized tokens in $U_{\text{top}}$ overlap with the top-ranked tokens in the token-level probability distribution. The adjusted generation distribution $P^{\prime}$ is computed as:
$P^{\prime}=P+\alpha H_{\text{norm}}(P)\cdot U_{\text{top}}$ (6)
where $P$ represents the original token-level probability distribution, and $\alpha H_{\text{norm}}(P)$ dynamically scales the adjustment based on model uncertainty.
5 Experiments
EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1
LLaMA2-7B Greedy 44.74 54.71 55.28 68.01 54.21 59.24 39.90 52.61 32.93 45.46 38.85 50.59 36.03 36.62 43.13 52.46
OURs 47.35 57.43 58.12 70.97 54.35 59.70 48.02 60.02 36.51 49.06 47.74 60.23 53.12 53.63 49.32 58.72
LLaMA2-7B -Chat Greedy 53.33 67.41 71.72 76.83 54.19 58.11 67.69 78.70 39.41 53.90 50.47 65.48 67.98 68.78 57.83 67.03
OURs 55.31 68.61 69.21 75.95 54.25 58.13 68.49 79.76 40.53 54.81 51.69 66.92 69.50 70.30 58.43 67.78
LLaMA2-13B Greedy 52.36 63.40 58.25 69.95 63.22 68.33 51.64 64.57 30.84 40.11 42.26 54.08 49.02 49.59 49.66 58.58
OURs 53.13 64.21 59.65 72.07 61.41 67.08 56.54 68.56 33.26 42.20 55.34 71.23 65.86 66.19 55.03 64.51
LLaMA2-13B -Chat Greedy 55.01 69.92 74.58 79.35 67.08 71.96 68.26 79.45 40.20 55.11 53.49 69.18 60.69 61.77 59.90 69.53
OURs 57.76 71.69 73.04 78.77 68.19 72.73 69.66 80.76 40.78 56.24 55.36 71.31 64.03 65.20 61.26 70.96
Mistral-7B Greedy 53.41 64.36 59.45 68.39 63.79 67.77 44.19 56.11 31.51 38.94 33.74 51.18 39.80 46.04 46.56 56.11
OURs 63.45 73.89 56.76 71.86 64.49 69.88 63.04 73.75 41.62 55.13 57.85 71.19 69.46 69.86 59.52 69.37
Mistral-7B -Instruct Greedy 58.70 72.18 69.64 75.61 44.42 49.63 67.28 79.37 39.79 54.72 52.29 66.93 66.90 67.83 57.00 66.61
OURs 60.55 73.49 69.70 75.96 47.17 52.65 68.30 80.62 40.85 56.23 54.78 69.72 71.38 72.12 58.96 68.68
Figure 5: Ablation Study for DAGCD. Left Part: Ablation 1 Detector Training Data Sizes, Center Part: Ablation 2 Top-Rank Constraint, Right Part: Ablation 3 Scaling Factor $\alpha$.
5.1 Experimental Setup
Datasets
We evaluate context faithfulness on Open-Book Question-Answering (QA) datasets, where each question is paired with external context containing the correct answer. This setup ensures that only context-grounded answers are considered valid, allowing for the assessment of whether the model generates context-faithful hallucinations based on answer accuracy. Specifically, We conducted experiments on seven open-book QA datasets. For further details, refer to Appendix D.1.
Implementation Details
We used 100 samples constructed from a single dataset (HotpotQA) as the training set to train the Context Utilization Detector. The scaling factor $\alpha$ is set to 2 for pre-trained models and 4 for instruction-tuned models to account for entropy variations. The logistic regression classifier and utilization distribution are computed using the top-10 attention heads, with adjustments restricted to the top-10 ranked tokens ($U_{\text{top}}$). Further detailed settings see Appendix D.2.
Metrics
Consistent with prior workJin et al. (2024a); Yuan et al. (2024); Wang et al. (2024), we use EM and F1 score metrics to evaluate the performance of the models on open-book QA datasets.
The LLMs used in our experiments and the baselines are detailed in Appendix D.3 and D.4.
5.2 Model Performance Comparison
Table 1 shows DAGCD’s effectiveness across diverse QA datasets and models. We also tested DAGCD on summarization tasks, yielding improvements. For details see Appendix D.5.
5.2.1 Dataset-Level Observations
HotpotQA, TriviaQA, SearchQA:
DAGCD excels in multi-paragraph reasoning and long-form retrieval tasks. It achieves the highest gains on HotpotQA with a 18.80% EM and 14.81% F1 improvement on Mistral-7B. DAGCD also outperforms baselines on TriviaQA and SearchQA, showing significant improvements across models.
SQuAD, NewsQA, NQ:
DAGCD demonstrates robust performance in single-paragraph and document-level tasks. On NQ, it achieves a 71.46% EM and 39.10% F1 improvement on Mistral-7B over greedy decoding, while delivering consistent gains across SQuAD and NewsQA datasets.
NQ-Swap:
In adversarial scenarios simulated by NQ-Swap, DAGCD shows notable improvements, including 74.52% EM and 51.74% F1 gains on Mistral-7B, highlighting its robustness.
5.2.2 Model-Level Observations
Model Families:
DAGCD enhances performance across LLaMA and Mistral families. On Mistral-7B, DAGCD improves EM by 27.86% on Mistral-7B and 14.33% on LLaMA2-7B compared to greedy decoding.
Model Sizes:
DAGCD improves performance across models of varying model sizes, achieving an average EM and F1 increase of 14.33% and 11.93% on LLaMA2-7B, and 10.82% and 10.12% on LLaMA2-13B, respectively.
Instruction-Tuned Models:
Instruction-tuned models, after fine-tuning, show reduced uncertainty in generation probabilities, limiting our method’s improvement margin. However, DAGCD still surpasses all baselines with the highest performance.
5.3 Ablation Study
We conducted three ablation experiments to evaluate the impact of variations in different modules on performance. Specifically, in each ablation experiment, the ablation module is adjusted while the other two modules are kept at their default settings.
Ablation 1: Detector Training Data Sizes
We tested the impact of detectors trained on different data sizes on actual inference performance. The results, shown in Figure 5 (left), demonstrate that our method consistently maintains strong performance across various training data sizes, achieving notable results even with just 100 training samples.
Ablation 2: Top-Rank Constraint
We evaluated various top-rank constraints on HotpotQA. Figure 5 (center), top-rank filtering reduces false positives, with F1 score initially improving as constraints loosen, then declining when overly relaxed.
Ablation 3: Scaling Factor $\alpha$
We evaluated the impact of different scaling factor $\alpha$ on model performance. The results, presented in Figure 5 (right), indicate that $\alpha$ determines the adjustment intensity applied to the original generation distribution. For pretrained models, optimal performance is achieved at $\alpha=2$, whereas for Chat models, the best performance is observed at $\alpha=4$.
Additional results and performance variations under different prompt templates see Appendix E.
6 Discussion and Analysis
Post-generation correction methods, such as CAD Shi et al. (2024b) and COIECD Yuan et al. (2024), improve contextual alignment but rely on multi-step processes, causing significant computational overhead. In contrast, DAGCD incorporates context adjustments during generation, ensuring both efficiency and real-time optimization.
Lightweight Context Utilization Detector Using a logistic regression-based Context Utilization Detector, our method enables real-time adjustments with minimal computational cost. This detector is more efficient than resource-heavy methods like integrated gradients or attention head manipulation.
Single-Pass Decoding with Real-Time Faithfulness Optimization By integrating the Context Utilization Detector directly into the decoding process, our method removes redundant steps like output comparisons or external consistency checks. During generation, attention-based context utilization signals are leveraged in real time to proactively enhance faithfulness. This single-step strategy ensures that the output aligns with the input context without extra post-processing, while maintaining the theoretical time complexity of greedy decoding.
By systematically analyzing attention mechanisms, our approach uncovers how retrieved context influences the generation process and provides interpretable insights into the behavior of the model.
Feature-Based Attention Analysis Using natural cases, such as failure in closed-book settings but success in open-book settings, we isolate attention patterns that are indicative of context utilization. A logistic regression classifier trained on these patterns identifies the relevant attention heads with high accuracy, quantifying their contributions.
Transparent Decision-Making The feature coefficients of the logistic regression model directly map to the importance of specific attention heads. This transparency allows for intuitive interpretation, clarifying which heads are most responsible for leveraging context tokens during generation.
7 Related Work
7.1 Context Faithfulness Hallucination
Current solutions to context faithfulness hallucination primarily focus on detection and mitigation. For detection, Lei et al. Lei et al. (2023) proposed a post-generation editing strategy using natural language inference to classify and revise hallucinated segments. Choi et al. Choi et al. (2023) introduced Knowledge-Constrained Decoding, detecting hallucinations during generation and reweighting token distributions to guide output. Chuang et al. Chuang et al. (2024a) proposed Lookback Lens Guided Decoding, selecting the most faithful output among candidates to improve consistency.
While effective, most existing methods face limitations such as high computational overhead and reliance on multiple decoding passes. In this work, we propose a real-time solution that integrates context utilization signals directly into the decoding process, achieving efficient and faithful generation, without the need for additional processing steps.
7.2 Attention and Interpretability
The attention mechanism provides valuable insights into how models prioritize different parts of an input sequence Clark et al. (2019); Geva et al. (2023) and has become central to understanding Transformer-based LLMs Vashishth et al. (2019); Hao et al. (2021); Zhao et al. (2024). In LLMs, attention heads often perform distinct roles, such as capturing syntactic dependencies or aligning semantic relationships Olsson et al. (2022b); Zheng et al. (2024); Jin et al. (2024b).
Unlike previous work, which focused on individual heads, our study examines the collaborative patterns of multiple attention heads. By analyzing how attention mechanisms collectively utilize contextual tokens, we provide a holistic view of their role in aligning outputs with user-provided context.
8 Conclusion
In this paper, we mitigate context faithfulness hallucinations in LLMs by proposing Dynamic Attention-Guided Context Decoding, a lightweight framework that integrates attention distributions and entropy-based uncertainty signals to amplify contextually relevant tokens during generation. Our analysis revealed a strong correlation between high uncertainty and hallucinations, and probing experiments validated that attention mechanisms encode signals indicative of contextual utilization, and further demonstrated that this signal is a fundamental mechanism in Transformer-based LLMs. Experiments across multiple open-book QA datasets demonstrated that DAGCD achieves consistent improvements in context faithfulness, robustness, and scalability, providing an effective solution for context-sensitive generation tasks.
Limitations
Dependency on Classifier Accuracy and Robustness to Noisy Contexts
Scaling Factor Adjustment for Model Characteristics
Generalization Across Tasks and Domains
Appendix A Details of Experiment "2.1 Uncertainty Leads to Unfaithful Answers"
A.1 Dataset Used for Analysis
We randomly sampled 6,000 instances from the MrQA training set Fisch et al. (2019), which consolidates six open-domain question answering datasets. Specifically, we selected 500 correctly answered and 500 incorrectly answered instances from each sub-dataset. An answer was considered correct if it achieved an Exact Match (EM) with the reference answer.
A.2 Model Used for Analysis, Prompts, and Decoding Method
We conduct our analysis using LLaMA2-7B Touvron et al. (2023) as the target model. The answers are generated using the following prompt:
"Given the following information: {context} Answer the following question based on the given information with one or a few words: {question} Answer:"
To ensure deterministic outputs, we employ greedy decoding.
A.3 Computation Process
For each sample, we calculated two metrics to quantify uncertainty and confidence. First, the Normalized Entropy (NE) measures the dispersion of probabilities across the vocabulary, providing an overall view of the model’s uncertainty. And NE is defined as: